OpenStudy (anonymous):

The radius of a circular puddle is growing at a rate of 25 cm/s. How fast is the area growing at the instant when it equals 81 cm?

OpenStudy (perl):

The area of a circular puddle is Pi * r^2

OpenStudy (anonymous):

so 81=pi(25)^2?

OpenStudy (perl):

The area of the puddle changes as the radius changes, but the radius changes as time changes or elapses.

OpenStudy (anonymous):

da/dt=pi*r^2? I'm not really sure how to set up the equation

OpenStudy (perl):

Area changes with radius, but radius changes with time

OpenStudy (perl):

thats a chain rule

OpenStudy (perl):

d/dt A(r(t)) = dA/dr * dr/dt

OpenStudy (perl):

or more simply we can write dA/dt = dA/dr * dr/dt

OpenStudy (perl):

let u = g(t) [f(u)] ' = f ' (u) * du/dt

OpenStudy (anonymous):

Okay, I'm not sure how to do it with the chain rule. I thought the setup would be something like dA/dt=2pi * r * dr/dt Would that be correct also?

OpenStudy (anonymous):

And then find dA/dt?

OpenStudy (perl):

yes that is correct